# Foundational Automatic Evaluators: Scaling Multi-Task Generative Evaluator Training for Reasoning-Centric Domains

### v1

@@ 4.2 Downstream evaluation
Reward model for inference-time scaling. We use the standardized setup in JETTS (Zhou et al., 2025b), which provides a set of 10 outputs from various generators and various benchmarks with corresponding correctness labels. Here, automatic evaluators rerank the responses, and performance is measured as the final performance of the evaluator-selected responses. We select the four most challenging benchmarks used in JETTS: MATH (Hendrycks et al., 2021), CHAMP (Mao et al., 2024), MBPP+ (Liu et al., 2023a), and BigCodeBench (Zhuo et al., 2024). We compare against strong judge models benchmarked previously on JETTS: SFR-Judge-8B,70B (Wang et al., 2024a), Skywork-Critic-8B,70B (Shiwen et al., 2024), Self-Taught-Evaluator-70B (Wang et al., 2024b), and CompassJudger-7B,14B (Cao et al., 2024). We utilize the default pairwise reranking setup and prompt FARE to produce a judgment directly without explanation, i.e., $y=(\emptyset,j)$.
@@
Verifer for GRPO training. We train with WebInstruct-Verified (Ma et al., 2025), a multi-domain reasoning dataset, covering math, chemistry, etc. Verifier impact is measured via the downstream performance of the trained policy model on a fixed evaluation suite of MMLU-Pro (Wang et al., 2024c), GPQA-Diamond (Rein et al., 2024), MATH-500, Minerva-Math (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024a), and AIME24 (Avg@32). We start from Qwen2.5-7B-Base (Yang et al., 2024) with the default reward setup as Ma et al. (2025) (see § C.2), and we compare against training with the string-matching verifier implemented by Ma et al. (2025).
Figure 4: Performance of downstream GRPO-trained model with FARE as reference-based verifiers. Moving from string based output matching to our models as generative verifiers brings tangible performance gains across both natural-language based (e.g., GPQA) and math settings.
Training with FARE-20B as a verifier improves downstream performance from 34.3 to 45.2, a nearly 11 point absolute gain, with improvements coming uniformly across the six benchmarks, as shown in Fig. 4. Notably, with 77% fewer gradient updates [^2] For the same group size, General-Reasoner-7B is trained for 700 steps with rollout batch size 768, whereas we train for 120 update steps with rollout batch size 1024. , our model was able to slightly beat the performance of General-Reasoner-7B (Ma et al., 2025) on several benchmarks: 38.9 vs. 38.8 on GPQA-Diamond, 38.4 vs 37.9 on OlympiadBench, and 13.9 vs. 13.8 on AIME24. This shows that using FARE-20B can significantly improve RL training convergence. Further, using FARE-8B and FARE-20B bring 8.8% and 14.1% relative gains over typically used string matching verifiers. These gains appear for both natural language (MMLU-Pro, GPQA) and math domains, showing that FARE can verify complex outputs across multiple challenging domains.
@@
Initialization for domain-specific continual finetuning. We continually finetune FARE-20B for code evaluation with one round of RS-SFT to produce FARE-20B-Code. We train with only 15K pairwise samples randomly chosen from AceCoder (Zeng et al., 2025). For evaluation, we use the recently released CodingJudgeBench (Jiang et al., 2025a), a pairwise benchmark covering code generation, code repair, and test-case quality evaluation tasks. Fig. 5, which reports consistent accuracy across the three splits from gpt-oss-20B, FARE-20B, FARE-20B-Code, and gpt-oss-120B for reference, shows that the first two splits are relatively easy, whereas test case evaluation is extremely difficult: gpt-oss-120B achieves only 32%. On the last task, FARE-20B improves over gpt-oss-20B by 11.67 absolute points, highlighting the benefits of large-scale evaluation training. Specialized continual training on coding tasks brings an additional 10.48 absolute point improvement, with FARE-20B-Code outperforming even gpt-oss-120B on average. In all, FARE can be readily adapted for specific applications with a small amount of domain-specific data.
@@ Appendix A Extended Related Work
Here, we cover recent advances in automatic evaluation that are non-generative. Scalar reward models (RMs) (Cobbe et al., 2021), which output a single scalar score for a given question $q$ and response $r$ were popularized originally within the context of reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Stiennon et al., 2020; Ziegler et al., 2019). Here, pairwise human preferences were used to train a RM, which is then applied during RL-based alignment, e.g., with PPO (Schulman et al., 2017) or DPO (Hosseini et al., 2024). Past work has focused on dataset curation (Jiang et al., 2023; Liu et al., 2024a; 2025a; Dong et al., 2024) and experimenting with different loss formulations (Liu et al., 2024a; Lou et al., 2024).
The above class of reward models operate at the outcome level, with the entire response being assigned a single score. Step-level reward models, known as process reward models, attempt to provide denser feedback by assigning each “step” in a response a score. Training PRMs relies on step-level labels from humans (Lightman et al., 2023) or models (Duan et al., 2025) or using Monte Carlo simulation to estimate step-level quality (Wang et al., 2023c; Luo et al., 2024; Xiong et al., 2025b; Zhang et al., 2025). Both approaches have associated drawbacks: Step-level annotation is rarely scalable, while Monte Carlo simulation requires careful filtering to ensure high quality.
While RMs and PRMs remain popular paradigms, recent approaches have found that generative evaluators can better leverage test-time compute for stronger evaluation performance (Zhang et al., 2024b; Mahan et al., 2024; Liu et al., 2025d), inspiring recent advances in generative evaluators, as discussed in § 2.
@@ B.1 Training data
We enumerate our training data sources in Table 5. We took efforts to decontaminate our training sets with N-gram matching approaches, following Guha et al. (2025). For the Synthetic data approach, we use 12 unique generators, covering a mixture of weak and strong models: Ministral-8B (Team, 2024), Mistral-Small 24B (Team, 2025a), Gemma 3 12B (Team et al., 2025), Qwen2.5 7B, 32B (Yang et al., 2024), Qwen-QwQ (Team, 2025b), Qwen3-30B-A3B (Yang et al., 2025), Llama-3.1-8B, Llama-3.3-70B (Dubey et al., 2024), GPT-4o (Hurst et al., 2024), and GPT-4.1-nano, 4.1 (OpenAI, 2025) for seed datasets with verifiable answers. To increase diversity we randomly select a prompt template from a preset list for each question in the seed dataset and sample multiple responses at varying temperatures (e.g., 0.0, 0.3, 0.5, 0.7,…). For open-ended datasets, such as tool-use datasets, we enumerated common errors found in tool calling outputs, such as invalid input types, missing arguments, malformed json format, etc., and then programmatically injected errors into ground-truth correct responses.
@@
ActPRM Duan et al. (2025) Step-Level Math Existing
Beavertails Preference Ji et al. (2023) Pairwise Safety Existing
@@
DeepScaleR Luo et al. (2025) Pairwise, Verification Math Synthetic
Folio Han et al. (2022) Pairwise, Verification NL Reasoning Synthetic
FoVer Kamoi et al. (2025) Step-Level Math Existing
HelpSteer Wang et al. (2023d) Single rating Chat Existing
HelpSteer2 Wang et al. (2024d) Single rating Chat Existing
HelpSteer3 Wang et al. (2025a; b) Pairwise Code, NL Reasoning Existing
HH-RLHF Harmless Bai et al. (2022); Ganguli et al. (2022) Pairwise Safety Existing
LAMP Chakrabarty et al. (2025) Pairwise, Single rating Chat Existing
MATH Hendrycks et al. (2021) Pairwise, Verification Math Synthetic
@@
OffsetBias Park et al. (2024) Pairwise Chat Existing
ReClor Yu et al. (2020) Pairwise, Verification NL Reasoning Synthetic
StepDPO Lai et al. (2024) Pairwise, Verification Math Existing
StrategyQA Geva et al. (2021) Pairwise, Verification NL Reasoning Synthetic
SWEGym Pan et al. (2024) Verification Code Existing
SWERank Reddy et al. (2025) Verification Code Existing
SynLogic Liu et al. (2025b) Pairwise, Verification NL Reasoning Synthetic
Tulu-V3-IF DPO data Lambert et al. (2024) Pairwise Chat Existing
@@
When2Call Preference Pairs Ross et al. (2025) Pairwise Tool-Use Synthetic
XLam-60K Liu et al. (2024c) Pairwise, Verification Tool-Use Synthetic
@@ B.2 Training details
We train batch size 128 and a constant learning rate of 1e-6 and choose per-iteration rollout batch sizes of 50,000 and 250,000 for FARE-8B and FARE-20B, respectively. In the latter case, we make the practical trade-off of RS-SFT iterations for training speed, reducing the number of times we need to reset the model for rollouts, etc. We use a modified version of the OpenRLHF framework (Hu et al., 2024a) for training.
@@ C.1 Core benchmarks
  • JudgeBench (Tan et al., 2024): A pairwise benchmark focused on evaluating LLM-as-judge models in reasoning settings, covering math, code, logical reasoning, and knowledge-based reasoning. Responses are generated using GPT-4o.
  • ReasoningJudgeBench (Xu et al., 2025b): A pairwise benchmark that covers more diverse reasoning settings, such as multi-hop, causal, and domain-specific reasoning. Responses are generated using GPT-4o.
  • PPE Correctness (Frick et al., 2024): A pairwise benchmark that covers reasoning and instruction following tasks with objectively correct answers, using seed datasets like MATH Hendrycks et al. (2021), GPQA (Rein et al., 2024), MBPP+ (Liu et al., 2023a), MMLU-Pro (Wang et al., 2024c), and IFEval (Zhou et al., 2023). Responses are generated using a variety of weaker models, e.g., Gemma-2-9B.
  • RM-Bench (Liu et al., 2024b): A pairwise benchmark that evaluates how robust evaluators are to stylistic biases by evaluating on pairs of responses with subtle yet critical differences.
  • When2Call (Ross et al., 2025): A pairwise benchmark that covers appropriate selection of tools (or refusals) in response to a user prompt. We use the LLM-as-judge test split, which comprises 300 unique prompts. Each prompt has four candidate answers (refusal, direct response, tool call, follow-up question), of which one response is correct. We form all pairs, yielding 900 total pairwise comparisons.
  • ProcessBench (Zheng et al., 2024): A step-level benchmark that evaluates the ability to identify step-level errors in mathematical reasoning across easy (GSM8K and MATH) and hard (Omni-Math and OlympiadBench) questions.
  • VerifyBench (Yan et al., 2025): A reference-based verification benchmark, comprised of Easy and Hard splits, that evaluates verifier ability to identify equivalent final answers.
@@
  • RISE-Judge (Yu et al., 2025a): Pairwise judges trained with SFT then DPO to perform pairwise evaluation.
  • Self-Taught Evaluators (Wang et al., 2024b): A pairwise judge trained with iterative SFT with training data generated in the loop.
  • EvalPlanner (Saha et al., 2025): Pairwise judges trained with iterative SFT and DPO on a small seed dataset, with an emphasis on learning how to plan for evaluation tasks.
  • RM-R1 (Chen et al., 2025b): A family of pairwise judges trained with GRPO, initialized from DeepSeek-distilled Qwen models.
  • J1 (Whitehouse et al., 2025): A pairwise and single-rating judge trained with GRPO.
  • CompassJudger (Cao et al., 2024): A family of foundational evaluators trained with large-scale SFT.
  • Atla Selene (Alexandru et al., 2025): A foundational evaluator trained with large-scale preference optimization.
We run gpt-oss variants with low reasoning, as (1) FARE-8B is trained initialized from gpt-oss-20B-low, and (2) evaluation often demands low-latency, making long CoT undesirable if they can be avoided. For GPT-5, we use the default API settings (medium reasoning). For pairwise benchmarks, we use reported scores from Whitehouse et al. (2025); Xu et al. (2025b); Liu et al. (2025d).
For ProcessBench, we use officially reported numbers in Zheng et al. (2024), which includes SkyworkPRMs (He et al., 2024b), Math Shepherd PRM (Wang et al., 2023c), ActPRM (Duan et al., 2025), and Qwen-Math PRMs (Zhang et al., 2025). We additionally report results from generative baselines: RL Tango (Zha et al., 2025) and StepWiser (Xiong et al., 2025b). For VerifyBench, we use reported scores from the original paper (Yan et al., 2025).
@@ C.2 Downstream settings
  • SFR-Judge-8B, 70B (Wang et al., 2024a): A family of multi-task evaluators. Among the highest performing small and large judges on JETTS.
  • Skywork-Critic-8B, 70B (Shiwen et al., 2024): Two pairwise-specfic evaluators that do not output explanations. Among the highest performing small and large judges on JETTS.
  • Self-Taught-Evaluator-70B (Wang et al., 2024b): A strong performing large judge model.
  • CompassJudger-7B, 14B (Cao et al., 2024): As described above.
Verifier during GRPO training. We adopt the settings of Ma et al. (2025), which train General-Reasoner, a family of reasoning LLMs of varying model sizes using the WebInstruct-Verified training dataset. In particular, we train with standard GRPO, i.e., without dynamic sampling or clip higher modifications, initializing from Qwen2.5-7B-Base. We the same conditional reward setup as General-Reasoner:
@@
The training framework is based on the verl framework (Sheng et al., 2024). We use rollout batch size 1024, max response length of 4096, group size of 8, a temperature of 1.0, a KL coefficient of 0.001, and a learning rate of 5e-7.
Initialization for domain-specific finetuning. We randomly sample 15,000 pairwise samples from AceCoder (Zeng et al., 2025) and perform one round of rejection sampling. We adopt training setup of FARE-20B: a direct judgment ratio of 60% and continuous curriculum. We train for 3 epochs with batch size 256 and cosine decay learning rate peaking at 1e-5. We then evaluate on CodingJudgeBench, reporting consistency accuracy.
Note that CodingJudgeBench reports an unconventional pairwise metric, employing Z-score normalization between the two consistency runs. Their implementation is not publicly available, and their paper lacks concrete implementation details. As such, we resort to consistent accuracy, which is more commonly used in pairwise benchmarks, e.g., (Tan et al., 2024; Li et al., 2023; Xu et al., 2025a; b).

### v2

@@ 4.2 Downstream evaluation
Reward model for inference-time scaling. We use the standardized setup in JETTS (zhou2025evaluating), which provides a set of 10 outputs from various generators and various benchmarks with corresponding correctness labels. Here, automatic evaluators rerank the responses, and performance is measured as the final performance of the evaluator-selected responses. We select the four most challenging benchmarks used in JETTS: MATH (hendrycks2021measuring), CHAMP (mao2024champ), MBPP+ (evalplus), and BigCodeBench (zhuo2024bigcodebench). We compare against strong judge models benchmarked previously on JETTS: SFR-Judge-8B,70B (wang2024direct), Skywork-Critic-8B,70B (skyworkcritic2024), Self-Taught-Evaluator-70B (wang2024self), and CompassJudger-7B,14B (cao2024compassjudger). We utilize the default pairwise reranking setup and prompt FARE to produce a judgment directly without explanation, i.e., $y=(\emptyset,j)$.
@@
Verifer for GRPO training. We train with WebInstruct-Verified (ma2025general), a multi-domain reasoning dataset, covering math, chemistry, etc. Verifier impact is measured via the downstream performance of the trained policy model on a fixed evaluation suite of MMLU-Pro (wang2024mmlu), GPQA-Diamond (rein2024gpqa), MATH-500, Minerva-Math (lewkowycz2022solving), OlympiadBench (he2024olympiadbench), and AIME24 (Avg@32). We start from Qwen2.5-7B-Base (yang2024qwen2) with the default reward setup as ma2025general (see § C.2), and we compare against training with the string-matching and trained verifier (General-Verifier) from ma2025general.
Figure 4: Performance of downstream GRPO-trained model with FARE as reference-based verifiers. Moving from string based output matching or weaker verifiers to FARE brings tangible performance gains across both natural-language based (e.g., GPQA) and math settings.
Training with FARE-20B as a verifier improves downstream performance from 34.3 to 45.2, a nearly 11 point absolute gain, with improvements coming uniformly across the six benchmarks, as shown in Fig. 4. Notably, with 77% fewer gradient updates [^2] For the same group size, General-Reasoner-7B is trained for 700 steps with rollout batch size 768, whereas we train for 120 update steps with rollout batch size 1024. , our model was able to slightly beat the performance of General-Reasoner-7B (ma2025general) on several benchmarks: 38.9 vs. 38.8 on GPQA-Diamond, 38.4 vs 37.9 on OlympiadBench, and 13.9 vs. 13.8 on AIME24. This shows that using FARE-20B can significantly improve RL training convergence. Further, using FARE-8B and FARE-20B bring 8.8% and 14.1% relative gains over typically used string matching verifiers and 4.4% and 9.4% over General-Verifier, which has been trained with in-training-distribution verification data. These gains appear for both natural language (MMLU-Pro, GPQA) and math domains, showing that FARE can verify complex outputs across multiple challenging domains.
@@
Initialization for domain-specific continual finetuning. We continually finetune FARE-20B for code evaluation with one round of RS-SFT to produce FARE-20B-Code. We train with only 15K pairwise samples randomly chosen from AceCoder (zeng2025acecoder). For evaluation, we use the recently released CodingJudgeBench (jiang2025codejudgebench), a pairwise benchmark covering code generation, code repair, and test-case quality evaluation tasks. Fig. 5, which reports consistent accuracy across the three splits from gpt-oss-20B, FARE-20B, FARE-20B-Code, and gpt-oss-120B for reference, shows that the first two splits are relatively easy, whereas test case evaluation is extremely difficult: gpt-oss-120B achieves only 32%. On the last task, FARE-20B improves over gpt-oss-20B by 11.67 absolute points, highlighting the benefits of large-scale evaluation training. Specialized continual training on coding tasks brings an additional 10.48 absolute point improvement, with FARE-20B-Code outperforming even gpt-oss-120B on average. In all, FARE can be readily adapted for specific applications with a small amount of domain-specific data.
@@ Appendix A Extended Related Work
Here, we cover recent advances in automatic evaluation that are non-generative. Scalar reward models (RMs) (cobbe2021training), which output a single scalar score for a given question $q$ and response $r$ were popularized originally within the context of reinforcement learning from human feedback (RLHF) (instructgpt_ouyang2022training; stiennon2020learning; ziegler2019fine). Here, pairwise human preferences were used to train a RM, which is then applied during RL-based alignment, e.g., with PPO (schulman2017proximal) or DPO (hosseini2024v). Past work has focused on dataset curation (jiang2023llm; liu2024skywork; liu2025skywork; dong2024rlhf) and experimenting with different loss formulations (liu2024skywork; lou2024uncertainty).
The above class of reward models operate at the outcome level, with the entire response being assigned a single score. Step-level reward models, known as process reward models, attempt to provide denser feedback by assigning each “step” in a response a score. Training PRMs relies on step-level labels from humans (lightman2023let) or models (duan2025efficient) or using Monte Carlo simulation to estimate step-level quality (wang2023math; luo2024improve; xiong2025stepwiser; prmlessons). Both approaches have associated drawbacks: Step-level annotation is rarely scalable, while Monte Carlo simulation requires careful filtering to ensure high quality.
While RMs and PRMs remain popular paradigms, recent approaches have found that generative evaluators can better leverage test-time compute for stronger evaluation performance (zhang2024generative; mahan2024generative; liu2025inference), inspiring recent advances in generative evaluators, as discussed in § 2.
@@ B.1 Training data
We enumerate our training data sources in Table 5. We took efforts to decontaminate our training sets with N-gram matching approaches, following guha2025openthoughts. For the Synthetic data approach, we use 12 unique generators, covering a mixture of weak and strong models: Ministral-8B (ministral8b), Mistral-Small 24B (mistral24b), Gemma 3 12B (team2025gemma), Qwen2.5 7B, 32B (yang2024qwen2), Qwen-QwQ (qwq32b), Qwen3-30B-A3B (yang2025qwen3), Llama-3.1-8B, Llama-3.3-70B (dubey2024llama), GPT-4o (hurst2024gpt), and GPT-4.1-nano, 4.1 (gpt41) for seed datasets with verifiable answers. To increase diversity we randomly select a prompt template from a preset list for each question in the seed dataset and sample multiple responses at varying temperatures (e.g., 0.0, 0.3, 0.5, 0.7,…). For open-ended datasets, such as tool-use datasets, we enumerated common errors found in tool calling outputs, such as invalid input types, missing arguments, malformed json format, etc., and then programmatically injected errors into ground-truth correct responses.
@@
ActPRM duan2025efficient Step-Level Math Existing
Beavertails Preference ji2023beavertails Pairwise Safety Existing
@@
DeepScaleR deepscaler2025 Pairwise, Verification Math Synthetic
Folio han2022folio Pairwise, Verification NL Reasoning Synthetic
FoVer kamoi2025training Step-Level Math Existing
HelpSteer wang2023helpsteer Single rating Chat Existing
HelpSteer2 wang2024helpsteer Single rating Chat Existing
HelpSteer3 wang2025helpsteer3; wang2025helpsteer3pref Pairwise Code, NL Reasoning Existing
HH-RLHF Harmless bai2022training; ganguli2022red Pairwise Safety Existing
LAMP chakrabarty2025can Pairwise, Single rating Chat Existing
MATH hendrycks2021measuring Pairwise, Verification Math Synthetic
@@
OffsetBias park2024offsetbias Pairwise Chat Existing
ReClor yu2020reclor Pairwise, Verification NL Reasoning Synthetic
StepDPO lai2024step Pairwise, Verification Math Existing
StrategyQA geva2021did Pairwise, Verification NL Reasoning Synthetic
SWEGym pan2024training Verification Code Existing
SWERank reddy2025swerank Verification Code Existing
SynLogic liu2025synlogic Pairwise, Verification NL Reasoning Synthetic
Tulu-V3-IF DPO data lambert2024t Pairwise Chat Existing
@@
When2Call Preference Pairs ross2025when2call Pairwise Tool-Use Synthetic
XLam-60K liu2024apigen Pairwise, Verification Tool-Use Synthetic
@@ B.2 Training details
We train batch size 128 and a constant learning rate of 1e-6 and choose per-iteration rollout batch sizes of 50,000 and 250,000 for FARE-8B and FARE-20B, respectively. In the latter case, we make the practical trade-off of RS-SFT iterations for training speed, reducing the number of times we need to reset the model for rollouts, etc. We use a modified version of the OpenRLHF framework (hu2024openrlhf) for training.
@@ C.1 Core benchmarks
  • JudgeBench (tan2024judgebench): A pairwise benchmark focused on evaluating LLM-as-judge models in reasoning settings, covering math, code, logical reasoning, and knowledge-based reasoning. Responses are generated using GPT-4o.
  • ReasoningJudgeBench (xu2025j4r): A pairwise benchmark that covers more diverse reasoning settings, such as multi-hop, causal, and domain-specific reasoning. Responses are generated using GPT-4o.
  • PPE Correctness (frick2024evaluate): A pairwise benchmark that covers reasoning and instruction following tasks with objectively correct answers, using seed datasets like MATH hendrycks2021measuring, GPQA (rein2024gpqa), MBPP+ (evalplus), MMLU-Pro (wang2024mmlu), and IFEval (zhou2023instruction). Responses are generated using a variety of weaker models, e.g., Gemma-2-9B.
  • RM-Bench (liu2024rm): A pairwise benchmark that evaluates how robust evaluators are to stylistic biases by evaluating on pairs of responses with subtle yet critical differences.
  • When2Call (ross2025when2call): A pairwise benchmark that covers appropriate selection of tools (or refusals) in response to a user prompt. We use the LLM-as-judge test split, which comprises 300 unique prompts. Each prompt has four candidate answers (refusal, direct response, tool call, follow-up question), of which one response is correct. We form all pairs, yielding 900 total pairwise comparisons.
  • ProcessBench (zheng2024processbench): A step-level benchmark that evaluates the ability to identify step-level errors in mathematical reasoning across easy (GSM8K and MATH) and hard (Omni-Math and OlympiadBench) questions.
  • VerifyBench (yan2025verifybench): A reference-based verification benchmark, comprised of Easy and Hard splits, that evaluates verifier ability to identify equivalent final answers.
@@
  • RISE-Judge (yu2025improve): Pairwise judges trained with SFT then DPO to perform pairwise evaluation.
  • Self-Taught Evaluators (wang2024self): A pairwise judge trained with iterative SFT with training data generated in the loop.
  • EvalPlanner (saha2025learning): Pairwise judges trained with iterative SFT and DPO on a small seed dataset, with an emphasis on learning how to plan for evaluation tasks.
  • RM-R1 (chen2025rmr1rewardmodelingreasoning): A family of pairwise judges trained with GRPO, initialized from DeepSeek-distilled Qwen models.
  • J1 (whitehouse2025j1): A pairwise and single-rating judge trained with GRPO.
  • CompassJudger (cao2024compassjudger): A family of foundational evaluators trained with large-scale SFT.
  • Atla Selene (alexandru2025atla): A foundational evaluator trained with large-scale preference optimization.
We run gpt-oss variants with low reasoning, as (1) FARE-8B is trained initialized from gpt-oss-20B-low, and (2) evaluation often demands low-latency, making long CoT undesirable if they can be avoided. For GPT-5, we use the default API settings (medium reasoning). For pairwise benchmarks, we use reported scores from whitehouse2025j1; xu2025j4r; liu2025inference.
For ProcessBench, we use officially reported numbers in zheng2024processbench, which includes SkyworkPRMs (he_2024_16998085), Math Shepherd PRM (wang2023math), ActPRM (duan2025efficient), and Qwen-Math PRMs (prmlessons). We additionally report results from generative baselines: RL Tango (zha2025rl) and StepWiser (xiong2025stepwiser). For VerifyBench, we use reported scores from the original paper (yan2025verifybench).
@@ C.2 Downstream settings
  • SFR-Judge-8B, 70B (wang2024direct): A family of multi-task evaluators. Among the highest performing small and large judges on JETTS.
  • Skywork-Critic-8B, 70B (skyworkcritic2024): Two pairwise-specfic evaluators that do not output explanations. Among the highest performing small and large judges on JETTS.
  • Self-Taught-Evaluator-70B (wang2024self): A strong performing large judge model.
  • CompassJudger-7B, 14B (cao2024compassjudger): As described above.
Verifier during GRPO training. We adopt the settings of ma2025general, which train General-Reasoner, a family of reasoning LLMs of varying model sizes using the WebInstruct-Verified training dataset. In particular, we train with standard GRPO, i.e., without dynamic sampling or clip higher modifications, initializing from Qwen2.5-7B-Base. We the same conditional reward setup as General-Reasoner:
@@
The training framework is based on the verl framework (sheng2024hybridflow). We use rollout batch size 1024, max response length of 4096, group size of 8, a temperature of 1.0, a KL coefficient of 0.001, and a learning rate of 5e-7.
Initialization for domain-specific finetuning. We randomly sample 15,000 pairwise samples from AceCoder (zeng2025acecoder) and perform one round of rejection sampling. We adopt training setup of FARE-20B: a direct judgment ratio of 60% and continuous curriculum. We train for 3 epochs with batch size 256 and cosine decay learning rate peaking at 1e-5. We then evaluate on CodingJudgeBench, reporting consistency accuracy.
Note that CodingJudgeBench reports an unconventional pairwise metric, employing Z-score normalization between the two consistency runs. Their implementation is not publicly available, and their paper lacks concrete implementation details. As such, we resort to consistent accuracy, which is more commonly used in pairwise benchmarks, e.g., (tan2024judgebench; li2023generative; xu2025does; xu2025j4r).